OpenStudy (anonymous):

Solve. x2 + 6x = 10 -3 ± sqrt 10 -3 ± sqrt 7 -3 ± sqrt 19 -3 ± sqrt 13

OpenStudy (anonymous):

...............this problem is wrong. or you typed it wrong. given this it would be 8x = 10. x = 10/8. is it x^2?

OpenStudy (anonymous):

yes x^2 + 6x = 10

OpenStudy (anonymous):

OK, so, we're solving for x, right? So you start by subtracting 10 from both sides getting you x^2 + 6x - 10 = 0. Are there any factors of 10 and 1 that can get you 6?

OpenStudy (anonymous):

no... i used the quadratic and i got it to -6 ± sqrt 76 ----------- 2

OpenStudy (anonymous):

hm. I see your problem.

OpenStudy (anonymous):

yep.. no idea what to do now

OpenStudy (anonymous):

convenient. simplify :). sqrt 76 = 2sqrt(19)

OpenStudy (anonymous):

you got the answer :)

OpenStudy (anonymous):

OK. Well, what about pulling out 4 (well, 2), and getting rid of it...or Ronncc did it...that was fast...

OpenStudy (anonymous):

anything even is tricky o_o